find the x intercept of the straight line passing through p=(2,3) and parallel to y+4x=3
Answer by nerdybill(7384)

find the x intercept of the straight line passing through p=(2,3) and parallel to y+4x=3
.
First, determine the equation of our new line.
Since it is parallel to
y+4x=3
it must have the same slope. Put into "slope-intercept" form:
y+4x=3
y = -4x+3
Therefore, slope (m) = -4
.
Using the slope and (2,3):
y-y1 = m(x-x1)
y-3 = -4(x-2)
y-3 = -4x+8
y = -4x+11 (equation of our new line)
.
To find the x-intercept, set y=0 and solve for x:
y = -4x+11
0 = -4x+11
-11 = -4x
11/4 = x
Therefore, the x-intercept is at (11/4, 0)